# Fisher-Geometric Diffusion in Stochastic Gradient Descent: Optimal Rates, Oracle Complexity, and Information-Theoretic Limits

### 10 Apr 2026 v3

@@ 8.1 Claim 1: The Lyapunov Equilibrium Law Is Quantitatively Predictive
with a two-dimensional sharp/flat curvature $H=\mathrm{diag}(1,0.1)$ and isotropic noise $F=I_{2}$, sweeping $b$ over a wide range. At stationarity, the covariance $\Sigma(b)$ solves the Lyapunov equation
$H\Sigma+\Sigma H\;=\;\frac{2}{b}\,F,$ (48)
so every stationary marginal variance scales as $1/b$; $\eta$ governs mixing time but not the fixed point. For the Euler–Maruyama simulation used here, the discrete Lyapunov equation (49) is the exact benchmark; the continuous equation (48) coincides at leading order because the stepsize satisfies $\eta\lambda_{\max}(H)\leq 0.05$.
Figure 2 overlays simulated stationary variances with the Lyapunov prediction and shows clean $1/b$ scaling across the batch sweep. The flat coordinate mixes more slowly (as expected under weak curvature) but converges to the same equilibrium law. This serves as a baseline check: the Fisher–Lyapunov equilibrium is quantitatively predictive, not merely qualitative. We observed early entry into the basin of Fisher geometry even from random starts; see Section 9.3 for a diagnostic based on the stability ratios $\delta_{F}$ and $\delta_{H}$.
Figure 2: OU benchmark: stationary variances vs. batch size $b$ (log scales). Solid: simulation; dashed: Lyapunov prediction from (48).
@@
8.1.2 Experiment 2: Anisotropic Fisher–Hessian Misalignment (Angle Sweep).
We keep the same curvature $H=\mathrm{diag}(1,0.1)$ but rotate an anisotropic noise-shape matrix relative to the curvature basis:
$F(\varphi)=R(\varphi)\begin{pmatrix}1.5&0\\[2.0pt] 0&0.5\end{pmatrix}R(\varphi)^{\top},\qquad\varphi\in[0,\pi/3],$
with fixed batch size, isolating alignment as an operational parameter. Because the simulator is Euler–Maruyama with finite time step, the appropriate comparator is the discrete Lyapunov equation: writing $A=I-\eta H\Delta t$ and $Q=(2\eta/b)F(\theta)\Delta t$, the stationary covariance solves
$\Sigma\;=\;A\Sigma A^{\top}+Q.$ (49)
Figure 3 shows a clear transfer of stationary variance across coordinates as $\varphi$ varies: when the noise eigenbasis rotates away from the curvature eigenbasis, equilibrium variance is redistributed in a predictable way, and the discrete Lyapunov benchmark tracks the simulation closely. The takeaway is structural: alignment governs which directions receive noise energy at equilibrium, even when scalar temperature $\eta/b$ is held fixed.
Figure 3: Angle sweep: stationary variances vs. misalignment angle. Bands show replicate variability; dashed curves are the discrete Lyapunov prediction from (49).
@@
8.1.3 Experiment 4: OU/Lyapunov Plateau in $d=10$ (Anisotropic vs. Isotropic).
Experiments 1–2 used $d=2$. We now move to $d=10$ to validate the Lyapunov equilibrium in a setting where anisotropy is richer. The curvature is $H=\mathrm{diag}(\lambda_{1},\ldots,\lambda_{10})$ with eigenvalues log-spaced between 1 and 100; the noise shape $G^{\star}$ is a random SPD matrix with eigenvalues in $[0.5,5]$, and we compare an isotropic control $G_{\mathrm{iso}}=\sigma^{2}I$ with $\sigma^{2}=\mathrm{Tr}(G^{\star})/d$ (trace-matched). Under constant step size $\eta=0.5/\lambda_{\max}(H)$ and batch size $b=64$, we run R=2, [^0] R=2{,}000 replicates and record the Fisher-metric risk $\mathbb{E}[\|\theta_{T}\|_{F^{\star}}^{2}]$ over increasing horizons $T\in\{300,800,2000,5000,10000\}$.
Table 4 confirms the Lyapunov prediction (Corollary 5.7): both the anisotropic and isotropic empirical risks converge to the predicted stationary level as $T$ grows, with convergence ratios $\widehat{R}/R_{\infty}$ approaching unity from both sides. The two plateaux nearly coincide because trace-matching is exact in this diagonal-$H$ setting; directional differences emerge only in higher-order structure (Claim 3 below).
@@
Table 4: OU/Lyapunov plateau ($d=10$): empirical Fisher-metric risk vs. discrete Lyapunov prediction $R_{\infty}=\mathrm{Tr}(F^{\star}\Sigma_{\infty})$. Convergence ratios $\widehat{R}_{T}/R_{\infty}\to 1$ confirm quantitative accuracy.
@@ 8.2 Claim 2: Decaying Step Sizes Recover a $1/N$ Rate with a Geometry-Determined Constant
The second claim is that under a Robbins–Monro step-size schedule the Fisher-metric risk decays as $C/N$ where $N=Tb$ is the total oracle budget and the constant $C$ is determined by the intrinsic geometry $(G^{\star},H^{\star},F^{\star})$. Specifically, the SA central-limit theorem gives an asymptotic covariance $V$ satisfying $({\eta_{0}H^{\star}-\tfrac{1}{2}I})V+V({\eta_{0}H^{\star}-\tfrac{1}{2}I})=\frac{\eta_{0}^{2}}{b}\,G^{\star}$, and $N\cdot\mathrm{Risk}\to b\,\mathrm{Tr}(F^{\star}V)$. This validates the matching upper and lower bounds (Theorem 6.5 and Proposition 6.10).
8.2.1 Experiment 5: $1/N$ Rate and Fisher Constant (Decaying Step Size).
We switch to a decaying step size $\eta_{t}=\eta_{0}/(t+t_{0})$ with $\eta_{0}=1.2$, $t_{0}=50$, and the same $(H,G^{\star},G_{\mathrm{iso}})$ from Experiment 4, with oracle budget $N=Tb$ and $b=64$. Figure 4 (left) plots the Fisher-metric risk on log–log axes: both anisotropic and isotropic risks decay as $1/N$ (fitted slopes $\approx-0.98$), confirming Theorem 6.5. The right panel shows the scaled risk $N\cdot\widehat{\mathrm{Risk}}$, which stabilizes near $\approx 14.2$ (anisotropic) and $\approx 14.0$ (isotropic)—both close to the SA–CLT prediction $b\,\mathrm{Tr}(F^{\star}V)$ where $V$ solves the Lyapunov equation above. The inset marks the auxiliary constant $\mathrm{Tr}(G^{\star}H^{-1})\approx 5.41$, which governs the constant-step OU stationary risk (Corollary 5.7); the ratio of the decaying-step plateau to this constant reflects the step-size schedule and the interplay between curvature and noise geometry.
Figure 4: Decaying step size, $d=10$. Left: Fisher-metric risk vs. NN (log–log); slopes $\approx-1$ confirm $1/N$ rate. Right: Scaled risk $N\cdot\widehat{\mathrm{Risk}}$ stabilizes; inset shows $C=\mathrm{Tr}(G^{\star}H^{-1})$.
@@
Table 5 summarizes the tightness of the bounds. The empirical risk is bounded between the information-theoretic lower bound (Proposition 6.10) and the upper bound (Theorem 6.5), with $N\cdot\text{Risk}$ stabilizing near $14.2$ (anisotropic) and $14.0$ (isotropic), both above the van Trees lower bound cd=10c\,d=10. At the largest budget (N=614,400N=614{,}400) the ratio of empirical constant to lower bound is approximately $1.4\times$, indicating near-tightness. The rate and constant are thus determined by $(G^{\star},H^{\star})$, so the oracle budget for a target risk can be computed from the problem’s intrinsic geometry.
Table 5: Upper vs. lower bound: scaled risk $N\cdot\widehat{\mathrm{Risk}}$ compared with the van Trees lower bound $cd=10$.
NN $N\!\cdot\!\widehat{\mathrm{Risk}}_{\mathrm{aniso}}$ $N\!\cdot\!\widehat{\mathrm{Risk}}_{\mathrm{iso}}$ $d=10$ (lower bound)
38 400 13.59 13.41 10
76 800 13.66 13.44 10
153 600 13.66 14.26 10
307 200 14.07 14.25 10
614 400 14.35 14.42 10
The SA–CLT Lyapunov prediction gives $b\,\mathrm{Tr}(F^{\star}V)\approx 14.2$ (anisotropic) and $\approx 14.0$ (isotropic); $N\cdot\mathrm{Risk}$ stabilizes above the van Trees lower bound cd=10c\,d=10 as expected.
8.3 Claim 3: Scalar Temperature Matching Fails—Directionality Is the Signature of Geometry
The third claim is that matching the total noise power (scalar temperature $\tau=\eta/b$) is insufficient: the directional distribution of noise energy is a distinguishing signature of the identified geometric model that no isotropic surrogate can reproduce. The Lyapunov trace identity $\mathrm{Tr}(H^{\star}\Sigma_{\infty})=\frac{\tau}{2}\mathrm{Tr}(G^{\star})$ depends only on $\mathrm{Tr}(G^{\star})$, so a trace-matched isotropic surrogate reproduces total risk by construction; but the directional distribution is entirely wrong, spreading residual error uniformly rather than concentrating it along top noise eigendirections.
8.3.1 Experiment 3: Geometry vs. Scalar Temperature (Rotation Diagnostic).
We compare a geometric OU model (diffusion covariance $\propto F_{\star}(\phi)$, rotating with angle $\phi$) against a trace-matched isotropic surrogate ($\sigma^{2}I$ with $\sigma^{2}=\mathrm{Tr}(F_{\star})/d$). At stationarity, the geometric covariance $\Sigma_{F}(\phi)$ generically rotates (producing nontrivial cross-covariance), while the isotropic $\Sigma_{I}(\phi)$ remains axis-aligned. Figures 57 confirm this: the ellipses tilt, cross-covariance $\Sigma_{12}$ tracks theory, and marginal variances redistribute—effects that temperature matching alone cannot capture.
Figure 5: Stationary covariance ellipses at three rotation angles: the identified equilibrium tilts with $\varphi$; the isotropic trace-matched surrogate remains axis-aligned.
$d=2$, $b=64$. Ellipses: 95% level sets of the Lyapunov solution.
Figure 6: Cross-covariance $\Sigma_{12}$ vs. rotation angle $\varphi$: isotropic trace-matched noise fails to reproduce off-diagonal covariance. Bands show $\pm 1$ s.d. across seeds.
$d=2$, 6 seeds, T=2,500T=2{,}500, burn-in=800{}=800.
Figure 7: Marginal variances $\Sigma_{11}$ (left) and $\Sigma_{22}$ (right) vs. angle: total noise energy may match, but directional allocation differs.
8.3.2 Experiment 6: Directional Variance Amplification.
@@ 8.3 Claim 3: Scalar Temperature Matching Fails—Directionality Is the Signature of Geometry
We now move to $d=10$ and ask whether anisotropy affects how risk is distributed across directions. For $k=1,2,3$ we project $\theta_{T}$ onto the top eigenvectors $v_{k}$ of $G^{\star}$ and compute the directional second moment $\mathbb{E}[\langle v_{k},\theta_{T}\rangle^{2}]$. Figure 8 (left) displays these on log–log axes: both models decay as $1/N$, but the anisotropic curve is systematically above the isotropic one along the top eigendirections. The right panel plots the ratio $\mathbb{E}_{\mathrm{aniso}}/\mathbb{E}_{\mathrm{iso}}$; at the largest budget the ratios are $1.65$, $1.45$, and $1.26$ for $k=1,2,3$, mirroring the eigenvalue ordering $\lambda_{1}>\lambda_{2}>\lambda_{3}$ of $G^{\star}$. These amplification ratios show that the identified geometry determines which parameter directions carry the most residual risk—information that scalar temperature models discard.
@@
Figure 8: Directional variance amplification ($d=10$). Left: Directional second moments along top eigenvectors of $G^{\star}$. Right: Ratio of anisotropic to isotropic directional variances; ratios >1>1 mirror the eigenvalue ordering of $G^{\star}$. Temperature matching preserves $\mathrm{Tr}(G^{\star})$ but not the eigen-structure; anisotropic noise concentrates residual risk along the top noise eigendirections.
Same as Experiment 5; top $k=1,2,3$ eigenvectors of $G^{\star}$.

### 15 Apr 2026 v4

@@ 8.2 Claim 2: The Lyapunov Geometry Is Visible in Low Dimensions
with $H=\mathrm{diag}(1,0.1)$ and varying noise shape $F$. This normalization is chosen for visualization only; it is not the literal fluctuation SDE of Theorem 5.7, which scales out the $b$-dependence into the $\sqrt{b/\eta}$ rescaled process $U^{\eta}$. At stationarity, the continuous-time Lyapunov equation for (54) reads $H\Sigma+\Sigma H=(2/b)F$, and the resulting marginal variances visualize the $1/b$ scaling, curvature–noise misalignment effects, and rotation diagnostics of the Lyapunov geometry in two dimensions.
Batch-size sweep ($1/b$ scaling).
With $F=I_{2}$ and a wide sweep of $b$, Figure 3 overlays simulated stationary variances with the fluctuation-scale Lyapunov prediction: clean $1/b$ scaling holds across the sweep, with the flat coordinate mixing more slowly (weaker curvature) but reaching the same equilibrium law.
Figure 3: Experiment 2, 2D batch-size sweep: stationary variances vs. $b$ (log scales); solid = simulation, dashed = fluctuation-scale Lyapunov prediction.
@@
Angle sweep (curvature–noise misalignment).
Rotating an anisotropic noise matrix $F(\varphi)=R(\varphi)\mathrm{diag}(1.5,0.5)R(\varphi)^{\top}$ relative to the curvature basis, Figure 4 shows a predictable transfer of stationary variance across coordinates as $\varphi$ varies: alignment governs which directions receive noise energy at equilibrium, even when the scalar temperature $\eta/b$ is held fixed. The discrete Lyapunov benchmark tracks the simulation closely.
Figure 4: Experiment 2, 2D angle sweep: stationary variances vs. misalignment angle $\varphi$. Bands show replicate variability; dashed curves are the discrete Lyapunov prediction.
@@
Rotation diagnostic (geometry vs. scalar temperature).
We compare a geometric OU with rotating diffusion $F_{\star}(\varphi)$ against a trace-matched isotropic surrogate $\sigma^{2}I$ with $\sigma^{2}=\mathrm{Tr}F_{\star}/d$. Figures 56 show that the geometric stationary covariance $\Sigma_{F}(\varphi)$ tilts with $\varphi$ (nonzero cross-covariance, redistributed marginals), while the isotropic surrogate remains axis-aligned at all angles. This is the signature the trace-matched surrogate cannot reproduce, and is the low-dimensional preview of the directional amplification measured in $d=10$ by Experiment 5.
Figure 5: Experiment 2, 2D rotation diagnostic: stationary covariance ellipses at three rotation angles. The geometric equilibrium tilts with $\varphi$; the isotropic trace-matched surrogate remains axis-aligned.
$d=2$, $b=64$; ellipses are 95% level sets of the Lyapunov solution.
Figure 6: Experiment 2, 2D rotation diagnostic: cross-covariance $\Sigma_{12}$ vs. rotation angle. Isotropic trace-matched noise fails to reproduce off-diagonal covariance; bands show $\pm 1$ s.d. across seeds.
$d=2$, 6 seeds, T=2,500T=2{,}500, burn-in =800=800.
8.3 Claim 3: The Lyapunov Plateau Holds Quantitatively in $d=10$
The 2D experiments display the qualitative geometry; we now move to $d=10$ for the first high-dimensional quantitative checkpoint. This is where the quantitative weight of the Lyapunov prediction is tested: Experiment 3 runs direct SGD on a $d=10$ problem and asks whether the empirically observed steady-state Fisher risk matches the level predicted by combining the identified covariance $G^{\star}$ with the discrete Lyapunov equation (Proposition 5.13). If the identification from Theorem 4.3 is quantitatively correct, both the anisotropic and isotropic empirical risks should converge to the Lyapunov-predicted plateau $\mathrm{Tr}(F^{\star}\Sigma_{\eta}^{\mathrm{lin}})$ as $T$ grows, with convergence ratios $\widehat{R}_{T}/R_{\infty}\to 1$.
8.3.1 Experiment 3: OU/Lyapunov Plateau in $d=10$ (Anisotropic vs. Isotropic).
The curvature is $H=\mathrm{diag}(\lambda_{1},\ldots,\lambda_{10})$ with eigenvalues log-spaced between 1 and 100, and we set $F^{\star}=H^{\star}$ throughout the synthetic $d=10$ benchmark (the quadratic-Gaussian surrogate in which curvature and Fisher information coincide, so the Fisher-metric risk $\mathbb{E}[\|\theta_{T}\|_{F^{\star}}^{2}]$ directly measures Hessian-weighted error). The noise shape $G^{\star}$ is a random SPD matrix with eigenvalues in $[0.5,5]$, and we compare an isotropic control $G_{\mathrm{iso}}=\sigma^{2}I$ with $\sigma^{2}=\mathrm{Tr}(G^{\star})/d$ (trace-matched). Under constant step size $\eta=0.5/\lambda_{\max}(H)$ and batch size $b=64$, we run R=2, [^0] R=2{,}000 replicates and record the Fisher-metric risk $\mathbb{E}[\|\theta_{T}\|_{F^{\star}}^{2}]$ over increasing horizons $T\in\{300,800,2000,5000,10000\}$.
Table 3 confirms the Lyapunov prediction: both the anisotropic and isotropic empirical risks converge to the predicted stationary level as $T$ grows, with convergence ratios $\widehat{R}_{T}/R_{\infty}$ within $1\%$ of unity across all horizons. The two plateaux nearly coincide because trace-matching is exact in this diagonal-$H$ setting; the distinction between anisotropic and isotropic geometry is therefore invisible at the level of scalar risk, and only emerges in higher-order structure (Experiment 5).
@@ 8.3 Claim 3: The Lyapunov Plateau Holds Quantitatively in $d=10$
Table 3: Experiment 3 (OU/Lyapunov plateau, $d=10$): empirical Fisher-metric risk vs. discrete Lyapunov prediction $R_{\infty}=\mathrm{Tr}(F^{\star}\Sigma_{\eta}^{\mathrm{lin}})$. Convergence ratios $\widehat{R}_{T}/R_{\infty}\to 1$ confirm quantitative accuracy.
@@ 8.4 Claim 4: Decaying Step Sizes Recover a $1/N$ Rate with a Geometry-Determined Constant
Under a Robbins–Monro step-size schedule the Fisher-metric risk decays as $C/N$ where $N=Tb$ is the total oracle budget and the constant $C$ is determined by the intrinsic geometry $(G^{\star},H^{\star},F^{\star})$. Specifically, the SA central-limit theorem gives an asymptotic covariance $V$ satisfying $({\eta_{0}H^{\star}-\tfrac{1}{2}I})V+V({\eta_{0}H^{\star}-\tfrac{1}{2}I})=\frac{\eta_{0}^{2}}{b}G^{\star}$, and $N\cdot\mathrm{Risk}\to b\,\mathrm{Tr}(F^{\star}V)$. This is consistent with the upper bound (Theorem 6.5) in rate order and with the i.i.d. parametric Fisher lower-bound benchmark (Proposition 6.8).
8.4.1 Experiment 4: $1/N$ Rate and Fisher Constant (Decaying Step Size).
We switch to a decaying step size $\eta_{t}=\eta_{0}/(t+t_{0})$ with $\eta_{0}=1.2$, $t_{0}=50$, and the same $(H,G^{\star},G_{\mathrm{iso}})$ from Experiment 3, with oracle budget $N=Tb$ and $b=64$. Figure 7 (left) plots the Fisher-metric risk on log–log axes: both anisotropic and isotropic risks decay as $1/N$ (fitted slopes $\approx-0.98$), consistent with Theorem 6.5 in rate order. The right panel shows the scaled risk $N\cdot\widehat{\mathrm{Risk}}$, which stabilizes near $\approx 14.2$ (anisotropic) and $\approx 14.0$ (isotropic)—both close to the SA–CLT prediction $b\,\mathrm{Tr}(F^{\star}V)$ where $V$ solves the Lyapunov equation above.
Figure 7: Experiment 4 (decaying step size, $d=10$). Left: Fisher-metric risk vs. NN (log–log); slopes $\approx-1$. Right: Scaled risk $N\cdot\widehat{\mathrm{Risk}}$ stabilizes; inset shows $C=\mathrm{Tr}(G^{\star}H^{-1})$.
@@
Interpretation: two related-but-distinct constants. The constant-step Lyapunov quantity $\mathrm{Tr}(G^{\star}H^{\star-1})$ (inset, $\approx 5.41$) and the decaying-step asymptotic constant $b\,\mathrm{Tr}(F^{\star}V)$ (plateau, $\approx 14.0$$14.2$) are related but not identical: the former is the stationary $\mathrm{Tr}(F^{\star}\Sigma_{\eta}^{\mathrm{lin}})$ of the constant-step discrete Lyapunov recursion at leading order in $\eta$, while the latter is the $N\to\infty$ asymptotic constant of the scaled risk under a Robbins–Monro schedule, which solves a different Lyapunov equation shifted by $\tfrac{1}{2}I$. The shift reflects the variance deflation introduced by the decaying schedule. Both constants are determined by the identified geometry $(G^{\star},H^{\star},F^{\star})$, but their numerical values differ, and the ratio between them depends on the schedule (here $\eta_{0}=1.2$, $t_{0}=50$). The empirical scaled risk $N\cdot\widehat{\mathrm{Risk}}$ lies above the i.i.d. parametric Fisher lower-bound benchmark of Proposition 6.8, consistent with rate-order matching; we do not claim a specific value for the theorem constant.
8.5 Claim 5: Directional Amplification Discriminates Geometry from Scalar Temperature
Experiments 3 and 4 compared anisotropic and isotropic noise at the level of trace-scale summaries, where trace-matching makes the two surrogates nearly indistinguishable. This is precisely where scalar-temperature reasoning fails. In the synthetic benchmark $F^{\star}=H^{\star}$, so the plotted Fisher-metric risk is $\mathrm{Tr}(F^{\star}\Sigma)=\mathrm{Tr}(H^{\star}\Sigma)$. Multiplying the fluctuation-scale Lyapunov equation $H^{\star}\Sigma_{U}+\Sigma_{U}H^{\star\top}=G^{\star}$ on the left by the identity and taking traces gives $2\,\mathrm{Tr}(H^{\star}\Sigma_{U})=\mathrm{Tr}(G^{\star})$, which depends on $G^{\star}$ only through its trace; a trace-matched isotropic surrogate ($G_{\mathrm{iso}}$ with $\mathrm{Tr}G_{\mathrm{iso}}=\mathrm{Tr}G^{\star}$) therefore reproduces the scalar Fisher-risk summary by construction, even though its directional distribution of residual risk is entirely different. The discriminating signature is directional: anisotropic noise concentrates stationary risk along the top eigendirections of $G^{\star}$, while the isotropic surrogate spreads residual error uniformly. Experiment 5 measures this direction-by-direction.
8.5.1 Experiment 5: Directional Variance Amplification.
@@ 8.5 Claim 5: Directional Amplification Discriminates Geometry from Scalar Temperature
For $k=1,2,3$ we project $\theta_{T}$ onto the top eigenvectors $v_{k}$ of $G^{\star}$ and compute the directional second moment $\mathbb{E}[\langle v_{k},\theta_{T}\rangle^{2}]$ under the same decaying-step setup as Experiment 4. Figure 8 (left) shows that both models decay as $1/N$, but the anisotropic curve is systematically above the isotropic one along the top eigendirections. The right panel plots the ratio $\mathbb{E}_{\mathrm{aniso}}/\mathbb{E}_{\mathrm{iso}}$: at the largest budget the empirical ratios are $(1.76,1.33,1.35)$ for $k=1,2,3$, close to the linearized Lyapunov prediction $(1.77,1.40,1.33)$ obtained from $H^{\star}\Sigma_{U}+\Sigma_{U}H^{\star\top}=G^{\star}$. The ratios are strictly greater than one in every direction and largest along the top eigenvector of $G^{\star}$. Trace-matched isotropic noise matches the scalar-risk summaries of Experiments 3–4 but cannot reproduce this direction-by-direction risk allocation; the identified geometry determines which parameter directions carry the most residual risk.
@@
Figure 8: Experiment 5 (directional variance amplification, $d=10$). Left: Directional second moments along top eigenvectors of $G^{\star}$. Right: Ratio of anisotropic to isotropic directional variances; ratios >1>1 mirror the eigenvalue ordering of $G^{\star}$. Trace-matched isotropic noise preserves $\mathrm{Tr}(G^{\star})$ but spreads residual risk uniformly; the identified anisotropic noise concentrates it along the top noise eigendirections.
Same setup as Experiment 4; top $k=1,2,3$ eigenvectors of $G^{\star}$.
8.6 Claim 6: Fluctuation-Scale $(\eta/b)$-Collapse on the Raw Iterate Path
Theorem 5.7 and Corollary 5.11 predict that, in the small-stepsize local regime, $\mathbb{E}[\|\theta_{T}-\theta^{\star}\|_{F^{\star}}^{2}]=(\eta/b)\,\mathrm{Tr}(F^{\star}\Sigma_{U})+o(\eta/b)$ with $\Sigma_{U}$ solving $H^{\star}\Sigma_{U}+\Sigma_{U}H^{\star\top}=G^{\star}(\theta^{\star})$. Equivalently, the scaled Fisher risk $(b/\eta)\,\widehat{\mathrm{Risk}}$ should collapse to a single $\eta,b$-free curve at leading order. This is the direct empirical signature of the fluctuation FCLT (Theorem 5.7) on the original iterate scale.
8.6.1 Experiment 6: Raw-Scale Collapse Diagnostic.
We sweep eight $(\eta,b)$ configurations with $\eta\in\{0.001,0.002,0.005\}$, $b\in\{32,64,128\}$ (so $\eta\lambda_{\max}(H^{\star})\in\{0.1,0.2,0.5\}$; $\eta/b\in[8\times 10^{-6},1.6\times 10^{-4}]$), and run direct SGD on the same $(H,F^{\star},G^{\star})$ as Experiments 3–4 for $T\in\{2000,5000,10000,20000\}$ with $R=600$ replicates per cell. Figure 9 plots $(b/\eta)\,\widehat{\mathbb{E}}[\|\theta_{T}\|_{F^{\star}}^{2}]$ against $T$ for each configuration. All curves collapse onto a common plateau near $\mathrm{Tr}(F^{\star}\Sigma_{U})\approx 10.22$ (black reference line), the fluctuation-scale Lyapunov prediction. The collapse across two orders of magnitude in $\eta/b$ is a direct verification of the $(\eta/b)$-scaling law on the raw iterate scale, independently of any continuous-time normalization convention.
Figure 9: Experiment 6 (fluctuation-scale collapse diagnostic; Theorem 5.7): scaled Fisher risk $(b/\eta)\,\widehat{\mathbb{E}}[\|\theta_{T}\|_{F^{\star}}^{2}]$ stabilizes at the fluctuation-scale Lyapunov prediction $\mathrm{Tr}(F^{\star}\Sigma_{U})$ (black line) across eight $(\eta,b)$ configurations.
$d=10$, direct SGD, $R=600$ replicates per cell; $\eta\in\{0.001,0.002,0.005\}$, $b\in\{32,64,128\}$.